OpenStudy (anonymous):

For exercise, Guy walks on a treadmill. The number of miles he walks can be shown with a graph of the equation y = 4x, where x represents hours. How far does Guy walk in 3 hours of exercise? A. 4 miles B. 8 miles C. 12 miles D. 14 miles

OpenStudy (anonymous):

C 12 miles

OpenStudy (nottim):

Nobody explains ANYTHING.

OpenStudy (anonymous):

ok but how do you find it

OpenStudy (anonymous):

The best way to find the answer is to plug in 3 to the equation and solve for Y. So, we know that y is the number of miles Guy will walk, right? And we know that x is the number of hours he's walking. Since y equals 4 times x, we have that y = 4 * 3 = 12.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (lilg132):

x represents hours equations is y = 4x so all you do is replace x with the number of hours which in the question is 3 hours y = 4(3) y = 12 dont forget to click good answer :)